mathslover (mathslover):

Hi help : \[\frac{T_{r+1}}{T_r}\] Binomial theorem

OpenStudy (anonymous):

Whats the quuestion?

mathslover (mathslover):

This was the actual question actually this was the topic : Greatest term in the expansion of (1+x)^n I want to calculate this : \[\frac{T_{r+1}}{T_r}\] for \[(1+x)^n\]

mathslover (mathslover):

i am getting this : \[\huge{\frac{x}{r[r*(n-r)]}}\]

OpenStudy (anonymous):

you should get \[\frac{n-r}{r+1}\]

OpenStudy (anonymous):

well multiplied by x ofcourse

OpenStudy (anonymous):